what does z equal to when the equation is 3z/10+3=-8
Answer by Alan3354(69443)

what does z equal to when the equation is 3z/10+3=-8
----------------
3z/10+3=-8
Subtract 3
3z/10 = -11
3z = -110
z = -110/3